# An $H^{1}$ multiplier theorem on anisotropic Hardy sapces

### v1

@@ Abstract.
A parallel result of Sledd–Stegenga’s $H^{1}\rightarrow L^{1}$ multiplier theorem was obtained on the $H^{1}$ space under the anisotropic settings. Based on the same technique, an $H^{1}\rightarrow L^{p}$ multiplier theorem is also proved for $1\leq p<\infty$.
@@ 1. Introduction
However, the anisotropic analog of Sledd–Stegenga’s criterion seems to be absent so far. We will close this gap. The statement of the criterion under the anisotropic settings is more complicated and requires more preparation. Moreover, some arguments in [8] does not work any more, and we need few modifications.
Also, there are problems in the original proofs of [8], which either omit many non-trivial steps or are too specific (only the 1-dimensional case is proved). Here we complete the missing details of the proof in [8].
@@ 2. Preliminaries
In this section, most of the text is adapted from the book [2].
@@ 2.1. Dilation matrix and its homogeneous norm
Let $A\in\mathbf{GL}(d,\mathbb{R})$ and $b\coloneqq|\det(A)|$. We say that $A$ is a dilation matrix, if all its eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$ have modulus strictly greater than $1$.
@@
The family of balls $\{\Delta_{k}\}_{k\in\mathbb{Z}}$ associated with $A$ around the origin is defined by
@@
Although all quasi-norms $\rho_{A}$ are equivalent (see [2], Chapter 1, Lemma 2.4), we will use the following specific one: From the nested property of the family $\{\Delta_{k}\}_{k\in\mathbb{Z}}$, we know that for every $x\in\mathbb{R}^{d}\setminus\{0\}$, there exists a unique $k\in\mathbb{Z}$ so that $x\in\Delta_{k+1}\setminus\Delta_{k}$. We define
@@ 2.2. Anisotropic $H^{1}$, and the atomic decomposition
Let $\varphi\in\mathcal{S}(\mathbb{R}^{d})$ with $\int\varphi\neq 0$. The anisotropic dilation of $\varphi$ is defined by $\varphi_{k}(\cdot)\coloneqq b^{k}\varphi(A^{k}\cdot)$. For $f\in L^{1}(\mathbb{R}^{d})$, the smooth maximal function of $f$ with respect to $\varphi$ is defined by
@@
The anisotropic $H^{1}$ space associated with the dilation $A$ is defined by
@@
The atomic decomposition of $H^{1}_{A}$ states that $f\in H^{1}_{A}$ if and only if there exists atoms $\{a_{i}\}_{i=1}^{\infty}$ and a sequence $(\lambda_{i})_{i=1}^{\infty}\in\ell^{1}$, so that $f=\sum_{i=1}^{\infty}\lambda_{i}a_{i}$. Moreover,
@@ 3. Main theorem and its proof
Let $\mu$ be a positive Borel measure on $\mathbb{R}^{d}\setminus\{0\}$. Then
@@ 3.1. Proof of the sufficient part
The region $(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})$ can be covered by $\mathcal{O}_{N}(1)$ translations of $(A^{\ast})^{l-N}(R^{\ast}(\alpha))$ with $\alpha\neq(0,0)$.
@@
where we used the hypothesis (3.2) in the last line.
We recall that $\Delta^{\ast}\subset R^{\ast}\subset(A^{\ast})^{N}(\Delta^{\ast})$. Applying the mapping $(A^{\ast})^{l-N}$ leads to
@@
Combining previous estimates, we can write
@@
The implicit constant depends (of course) on $N,b$ and $\zeta_{-}$, but not on $k$.
@@
In the original paper [8], the estimate over $(A^{\ast})^{-(k+N)}(R^{\ast})$ is omitted and the authors said it was an easy consequence of (3.2). This is not quite accurate. In [8], the authors write the summation (3.2) over all $\alpha\in\mathbb{Z}^{d}$, which trivially implies the estimate over $(A^{\ast})^{-(k+N)}(R^{\ast})$. However, the correct condition is the sum over $\alpha\in\mathbb{Z}^{d}\setminus\{0\}$. In this case, we need few more work.
Having completed the estimate of the integral on $(A^{\ast})^{-(k+N)}(R^{\ast})$, we now turn to control the integral on the complement $\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})$, it is easy to check that $\mathbb{R}^{d}\setminus R^{\ast}=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}R^{\ast}(\alpha)$, applying the invertible mapping $(A^{\ast})^{-(k+N)}$ gives
$\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha)),$
and the union is a disjoint union. Therefore
$\int_{\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}|\,\mathrm{d}\mu$ $=\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\int_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|\,\mathrm{d}\mu$
@@
$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\mu\Big{(}(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))\Big{)}^{2}\lesssim 1.$
@@
(3.3) $\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}\lesssim 1.$
@@
$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}=b^{2(k+N)}\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\sup_{\eta\in R^{\ast}(\alpha)}\Big{|}\mathcal{F}\big{(}a\circ A^{k+N}\big{)}(\eta)\Big{|}^{2}$
@@
In [8], the authors wrote that the higher-dimensional case is somewhat complicated so the proof was omitted. We give a proof of the two-dimensional case. For general $d\geq 3$, the proof below also works, but the notations will become more annoyed.
@@
Let us first see how to control $|b_{Q}|^{2}$, which is easier. The Cauchy–Schwarz inequality implies $|b_{Q}|^{2}\leq\frac{1}{|Q|}\int_{Q}|f|^{2}$. Summing over all $\alpha\in\mathbb{Z}^{2}$, we obtain
@@
$|f(x)-b_{Q}|=$ $\Big{|}f(x)-\frac{1}{|Q|}\int_{Q}f(y)\ \mathrm{d}y\Big{|}\leq\frac{1}{|Q|}\int_{Q}|f(x)-f(y)|\ \mathrm{d}y$
$=$ $\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}|f(x_{1},x_{2})-f(y_{1},y_{2})|\ \mathrm{d}y_{2}\mathrm{d}y_{1}$
$\leq$ $\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}|f(x_{1},x_{2})-f(y_{1},x_{2})|\ \mathrm{d}y_{2}\mathrm{d}y_{1}$
$+$ $\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}|f(y_{1},x_{2})-f(y_{1},y_{2})|\ \mathrm{d}y_{2}\mathrm{d}y_{1}.$
Now we introduce the notations
@@
The integrand $\partial_{1}f(t,x_{2})$ is independent of $y_{1}$, so by Cauchy–Schwarz,
@@
In [8], it was proved that for any interval $I$ and and smooth $f:I\rightarrow\mathbb{R}$,
@@
By hypothesis, the atom $a$ is supported in $\Delta_{k}=A^{k}(\Delta_{0})$, so the function $a\circ A^{k+N}$ is supported in $A^{-N}(\Delta_{0})$. Using $\partial^{\alpha}\widehat{f}(\xi)=\mathcal{F}\big{(}(-2\pi\mathrm{i}x)^{\alpha}f(x)\big{)}(\xi)$ and the Plancherel’s identity,
@@
where the implicit constant may depends on $b,N,\Delta_{0}$ and so on, but independent of $k$. Other terms can be treated by the same way. Similar argument gives the higher dimensional result of (3.3) for $d\geq 3$.
@@ 3.2. Proof of the necessary part
We are going to use the following well-known result on Bochner–Riesz multiplier.
@@
Suppose that $\lambda>-1$. The function $m_{\lambda}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined by
@@
This result is standard and can be found in many textbook in Fourier analysis, for example, at the Appendix B of [5]. We remind the reader that $\Delta^{\ast}_{0}$ is the ellipse associated with the matrix $A^{\ast}$, so we may assume that $\Delta^{\ast}_{0}=P^{-1}(B(0,1))$ for certain $P\in\mathbf{GL}(d,\mathbb{R})$. Moreover, after rescaling, we can assume that $|\Delta_{0}^{\ast}|\approx 1$.
@@
The argument here is adapted from [4]. We recall that $\Delta^{\ast}_{0}=\{x\in\mathbb{R}^{d}:|x|_{\ast}\leq 1\}=\{x\in\mathbb{R}^{d}:|Px|\leq 1\}$, hence $\mathrm{supp}(\chi_{k,\lambda})\subset\Delta^{\ast}_{k}$.
We begin by showing that $\int f=0$. This is easy. In fact, we have
@@
$+\sum_{l=0}^{\infty}\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
For the integration on $A^{-k}(\Delta_{0})$, the trivial fact $|A^{k}x|\geq 0$ implies that
@@
We decompose, for $l\geq 0$, the integral $\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}$ as
@@
For $x\in A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})$, we have
$A^{k}(x)\in A^{l+1}(\Delta_{0})\setminus A^{l}(\Delta_{0}),\text{ and }\rho(A^{k}x)=b^{l}\geq 1.$
@@
$\lesssim\frac{1}{b^{l(d+1+2\lambda)\zeta_{-}}}\int_{A^{l+1-k}(\Delta_{0})}|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}$
@@
Therefore, by telescoping,
$|\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|\lesssim(l+1)b\|u\|_{\mathrm{BMO}},$
and
@@
Summing over $l$ gives
@@
Now we back to the proof of the necessary part. Suppose that (3.1) holds. Applying (3.1) for $f=g\cdot\widehat{\chi_{k,\lambda}}$, and $\|g\|_{2}\leq 1$,
$\Big{|}\int\widehat{g}\ast\chi_{k,\lambda}\,\mathrm{d}\mu\Big{|}=\int|\widehat{f}|\,\mathrm{d}\mu\lesssim\|f\|_{H^{1}}\lesssim b^{\frac{k}{2}}.$
Notice that
$\int\left(\widehat{g}\ast\chi_{k,\lambda}\right)(x)\,\mathrm{d}\mu(x)=\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y.$
Therefore
$\Big{|}\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y\Big{|}\lesssim b^{\frac{k}{2}}.$
Recall that by our hypothesis, the Fourier transform $\widehat{g}$ is supported in $(\Delta^{\ast}_{k})^{c}$, so by duality,
@@
In consequence, the norm $|P(A^{\ast})^{-k}x|\leq 1-\epsilon$, and $\chi_{k,\lambda}(x)\gtrsim_{\epsilon,\lambda}1$.
Therefore, we have
@@
$\gtrsim_{\epsilon,\lambda}\mu\big{(}y+(1-\epsilon)\Delta^{\ast}_{k}\big{)}$
@@
We already knew that $(A^{\ast})^{-N}(R^{\ast})\subset\Delta^{\ast}_{0}$, so
@@
Replacing $(1-\epsilon)\Delta^{\ast}_{k}$ by a smaller subset $(A^{\ast})^{k}(cR^{\ast})$ leads to
@@
For $\alpha$ fixed, we decompose $R^{\ast}(\alpha)$ into 9 congruent parts $r_{i}$. It is evident that, the translated rectangle $y+\frac{1}{2}R^{\ast}$ contains $r_{i}$, whenever $y$ is near the center of $r_{i}$. Thus we may define a smaller rectangle concentric with $r_{i}$:
@@
A trivial verification shows that $|r_{i}|/|r_{i}^{\prime}|\approx 1$. For general $c>0$, this area quotient depends of course on $c$, but the important point is that it is independent of $k$.
@@
We recall that $|r^{\prime}_{i}|\approx|r_{i}|=|R^{\ast}|/9$, and $|(A^{\ast})^{k}(r^{\prime}_{i})|=b^{k}|r^{\prime}_{i}|$. Thus the above inequality can be further written as
@@
Now back to (3.4), the following inequality holds:

### v3

# An $H^{1}$ multiplier theorem on anisotropic Hardy spaces

@@ Abstract.
We obtain a parallel result of Sledd–Stegenga’s $H^{1}\rightarrow L^{1}$ multiplier theorem under the anisotropic settings. Based on the same technique, we also prove an $H^{1}\rightarrow L^{p}$ multiplier theorem for $1\leq p<\infty$.
@@ 1. Introduction
However, the anisotropic analog of Sledd–Stegenga’s criterion seems to be absent so far. We close this gap in this note. The statement of the criterion under the anisotropic settings is more complicated and requires more preparation. Moreover, some arguments in [8] does not work any more, and we need few modifications.
Also, there are gaps in the original proofs of [8], which either omit many non-trivial steps or are too specific (only the 1-dimensional case is proved, whereas the higher-dimensional case is not a direct result thereof). Here we complete the missing details in [8] incidentally.
2. Acknowledgement
This paper was completed at Nicolaus Copernicus University in Toruń, where the author was a postdoctoral researcher at that time. The author would like to thank Yuri Tomilov for introducing him to the paper [8]. The author also express the gratitude to Marcing Bownik for introducing him the theory of anisotropic Hardy spaces.
@@ 3. Preliminaries
In this section, most of the text is adapted from [3].
@@ 3.1. Dilation matrix and its homogeneous norm
A matrix $A\in\mathbf{GL}(d,\mathbb{R})$ is a dilation matrix, if all its eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$ have modulus strictly greater than $1$.
@@
The family of balls $\{\Delta_{k}\}_{k\in\mathbb{Z}}$ associated with AA around the origin is
@@
Although all quasi-norms $\rho_{A}$ are equivalent (see [3], Chapter 1, Lemma 2.4), we will use the following specific one: From the nested property of the family $\{\Delta_{k}\}_{k\in\mathbb{Z}}$, for every $x\in\mathbb{R}^{d}\setminus\{0\}$, there exists unique $k\in\mathbb{Z}$ so that $x\in\Delta_{k+1}\setminus\Delta_{k}$. We define
@@ 3.2. Anisotropic $H^{1}$, and the atomic decomposition
Let $\varphi\in\mathcal{S}(\mathbb{R}^{d})$ with $\int\varphi\neq 0$. The anisotropic dilation of $\varphi$ is defined by $\varphi_{k}(\cdot)\coloneqq b^{k}\varphi(A^{k}\cdot)$. For $f\in L^{1}(\mathbb{R}^{d})$, the smooth maximal function of $f$ with respect to $\varphi$ is
@@
The anisotropic $H^{1}$ space associated with the dilation $A$ is
@@
The atomic decomposition of $H^{1}_{A}$ states that $f\in H^{1}_{A}$ if and only if there exist atoms $\{a_{i}\}_{i=1}^{\infty}$ and a sequence $(\lambda_{i})_{i=1}^{\infty}\in\ell^{1}$, so that $f=\sum_{i=1}^{\infty}\lambda_{i}a_{i}$. Moreover,
@@ 4. Main theorem and its proof
Let $\mu$ be a positive Borel measure on $\mathbb{R}^{d}\setminus\{0\}$, then
@@ 4.1. Proof of the sufficient part
The region $(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})$ can be covered by $\mathcal{O}_{N}(1)$ translations of $(A^{\ast})^{l-N}(R^{\ast}(\alpha))$ with $\alpha\neq 0$.
@@
where we used the hypothesis (4.2) in the last inequality.
Since $\Delta^{\ast}\subset R^{\ast}\subset(A^{\ast})^{N}(\Delta^{\ast})$, applying the mapping $(A^{\ast})^{l-N}$ leads to
@@
Combining previous estimates,
@@
The implicit constants are independent of $k$.
@@
In the original paper [8], the estimate over $(A^{\ast})^{-(k+N)}(R^{\ast})$ is omitted, and the authors said it was an easy consequence of (4.2). This is not quite accurate. In [8], the authors write the summation (4.2) over all $\alpha\in\mathbb{Z}^{d}$, which trivially implies the estimate over $(A^{\ast})^{-(k+N)}(R^{\ast})$. However, the correct condition is the sum over $\alpha\in\mathbb{Z}^{d}\setminus\{0\}$.
We now turn to control the integral on the complement $\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})$, it is easy to check that $\mathbb{R}^{d}\setminus R^{\ast}=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}R^{\ast}(\alpha)$, applying the invertible mapping $(A^{\ast})^{-(k+N)}$ gives
$\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha)),$
and the union is a disjoint union. Therefore,
$\int_{\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}|\,\mathrm{d}\mu$ $=\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\int_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|\,\mathrm{d}\mu$
@@
$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\mu\Big((A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))\Big)^{2}\lesssim 1.$
@@
(4.3) $\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}\lesssim 1.$
@@
$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}=b^{2(k+N)}\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\sup_{\eta\in R^{\ast}(\alpha)}\Big|\mathcal{F}\big(a\circ A^{k+N}\big)(\eta)\Big|^{2}$
@@
In [8], the authors wrote that the higher-dimensional case is somewhat complicated, so the proof was omitted. We give a proof of the two-dimensional case. For general $d\geq 3$, the proof below also works, but the notations will become more annoyed.
@@
Control $|b_{Q}|^{2}$: The Cauchy–Schwarz inequality implies $|b_{Q}|^{2}\leq\frac{1}{|Q|}\int_{Q}|f|^{2}$. Summing over all $\alpha\in\mathbb{Z}^{2}$, we obtain
@@
[…]
where we introduce the notations
@@
The integrand $\partial_{1}f(t,x_{2})$ is independent of $y_{1}$, by Cauchy–Schwarz,
@@
In [8], it was proved that for any interval $I$ and smooth $f:I\rightarrow\mathbb{R}$,
@@
By hypothesis, the atom $a$ is supported in $\Delta_{k}=A^{k}(\Delta_{0})$, so the function $a\circ A^{k+N}$ is supported in $A^{-N}(\Delta_{0})$ and bounded by $|A^{-N}(\Delta_{0})|^{-1}$. Using $\partial^{\alpha}\widehat{f}(\xi)=\mathcal{F}\big((-2\pi\mathrm{i}x)^{\alpha}f(x)\big)(\xi)$ and the Plancherel’s identity,
@@
where implicit constants may depend on $b,N,\Delta_{0}$ and so on, but are independent of $k$. Other terms can be treated by the same way. Similar argument gives the higher dimensional result of (4.3) for $d\geq 3$.
@@ 4.2. Proof of the necessary part
We are going to use a well-known result on Bochner–Riesz multiplier. Suppose that $\lambda>-1$. The function $m_{\lambda}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined by
@@
This result is standard and can be found in many textbooks in Fourier analysis, for example, at the Appendix B of [5].
@@
The argument here is adapted from [4]. We remind the reader that $\Delta^{\ast}_{0}$ is the ellipse associated with the matrix $A^{\ast}$, so we may assume that $\Delta^{\ast}_{0}=P^{-1}(B(0,1))$ for certain $P\in\mathbf{GL}(d,\mathbb{R})$. Moreover, after re-scaling, we can assume that $|\Delta_{0}^{\ast}|\approx 1$. It is trivial that $\mathrm{supp}(\chi_{k,\lambda})\subset\Delta^{\ast}_{k}$.
We begin by showing that $\int f=0$. Using the support conditions of $\widehat{g}$ and $\chi_{k,\lambda}$,
@@
$+\sum_{l=0}^{\infty}\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}.$
The trivial fact $|A^{k}x|\geq 0$ implies that
@@
We decompose, for each $l\geq 0$, the integral $\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}$ as
@@
For $x\in A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})$,
$A^{k}x\in A^{l+1}(\Delta_{0})\setminus A^{l}(\Delta_{0}),\text{ and }\rho(A^{k}x)=b^{l}\geq 1.$
@@
$\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$ $\lesssim\int_{A^{l+1-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}}{b^{l(d+1+2\lambda)\zeta_{-}}}$
@@
By telescoping,
$|\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|\lesssim(l+1)b\|u\|_{\mathrm{BMO}}.$
Consequently,
@@
Summing over $l\in\mathbb{N}$,
@@
Back to the proof of the necessary part. Suppose that (4.1) holds. Applying (4.1) for $f=g\cdot\widehat{\chi_{k,\lambda}}$ with $\|g\|_{2}\leq 1$,
$\Big|\int\widehat{g}\ast\chi_{k,\lambda}\,\mathrm{d}\mu\Big|=\int|\widehat{f}|\,\mathrm{d}\mu\lesssim\|f\|_{H^{1}}\lesssim b^{\frac{k}{2}},$
while
$\int\widehat{g}\ast\chi_{k,\lambda}\,\mathrm{d}\mu=\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y.$
Therefore,
$\Big|\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y\Big|\lesssim b^{\frac{k}{2}}$
By our hypothesis, the Fourier transform $\widehat{g}$ is supported in $(\Delta^{\ast}_{k})^{c}$, so by duality,
@@
In consequence, the norm $|P(A^{\ast})^{-k}x|\leq 1-\epsilon$, and $\chi_{k,\lambda}(x)\gtrsim_{\epsilon,\lambda}1$. Thus,
@@
$\gtrsim_{\epsilon,\lambda}\mu\big(y+(1-\epsilon)\Delta^{\ast}_{k}\big).$
@@
Also, since $(A^{\ast})^{-N}(R^{\ast})\subset\Delta^{\ast}_{0}$,
@@
Replacing $(1-\epsilon)\Delta^{\ast}_{k}$ by the smaller subset $(A^{\ast})^{k}(cR^{\ast})$ leads to
@@
For $\alpha$ fixed, we decompose $R^{\ast}(\alpha)$ into 9 congruent parts $r_{i}$. It is evident that, the translated rectangle $y+\frac{1}{2}R^{\ast}$ contains $r_{i}$, whenever $y$ is near the center of $r_{i}$. Thus, we define a smaller rectangle concentric with $r_{i}$:
@@
A trivial verification shows that $|r_{i}|/|r_{i}^{\prime}|\approx 1$. For general $c>0$, the number of congruent parts into which $R^{\ast}(\alpha)$ is divided, and the area quotient $|r_{i}|/|r_{i}^{\prime}|$ depend of course on $c$, but the important point is that they are independent of $k$.
@@
Since $|r^{\prime}_{i}|\approx|r_{i}|=|R^{\ast}|/9$, and $|(A^{\ast})^{k}(r^{\prime}_{i})|=b^{k}|r^{\prime}_{i}|$, the above inequality can be further written as
@@
Now back to (4.4), the following inequality holds,